OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

add 2 equations left hand side and right hand side seperately the y terms will get cancelled

Parth (parthkohli):

Add both equations(elimination).

OpenStudy (unklerhaukus):

add the equations together, the y-s will cancel

Parth (parthkohli):

Is that enough for you to manage it? @rebecca1233

OpenStudy (unklerhaukus):

you can solve it in your head this way

Parth (parthkohli):

And by adding both equations, we mean adding the LHS's and RHS's.

OpenStudy (anonymous):

can you help me work it out please

OpenStudy (unklerhaukus):

(x - y)+ (2x + y) = (13) + (2) x - y+ 2x + y = 13 + 2 (x + 2x) +( y - y) = (13 + 2)

OpenStudy (anonymous):

alright & which method is this cause i was gnna do substitution

OpenStudy (anonymous):

x=5

OpenStudy (unklerhaukus):

if you wanted to use substitution i would target the y in the second equation

OpenStudy (anonymous):

ok what method did u use

OpenStudy (unklerhaukus):

Parth says its called (elimination)

OpenStudy (anonymous):

ok.. and i got y = -8

OpenStudy (unklerhaukus):

which method did you use, correct by the way

OpenStudy (anonymous):

yes well done x=5 y=-8

OpenStudy (anonymous):

Thanks.. can you guys help with another question i post

OpenStudy (unklerhaukus):

x - y = 13 2x + y = 2 [1 -1 | x| 13 ] [2 1 | y | 2 ] [1 -1 | x| 13 ] [0 3 | y | -24 ] R_2 - 2R_1 [1 -1 | x | 13 ] [0 1 | y | -8 ] R_2 / 3 [1 0 | x | 5 ] R_1+R_2 [0 1 | y | -8 ] x=5 y=-8